A particle leaves the 'origin with velocity u. If
it moves with an acceleration -Liv being
v2
the velocity at any instant, show that the
distance x travelled in time t is given by
4 px= (3 lit+ 112)4/3 - u4.
1
Expert's answer
2020-10-09T07:13:53-0400

As per the given question,

It is given that acceleration of the particle "(a)=\frac{Li}{v^2}"

We know that,

"a=\frac{dv}{dt}"

"\Rightarrow dv= adt"

Now, substituting the value of a,

"\int_u^v dv=\int_0^t\frac{Li}{v^2}dt"


"\Rightarrow \int_u^vv^2dv=\int_0^tLidt"

"\Rightarrow [\frac{v^3}{3}]_u^v=Lit"

"\Rightarrow v^3=3Lit+u^3"


"\Rightarrow v=(Lit+u^3)^{1/3}"

We know that,

"v=\frac{dx}{dt}"

"\Rightarrow dx=vdt"

Integrating both side with respect to t,

"x=\int_0^t (Lit+u^3)^{1/3}dt"


"\Rightarrow x =[\frac{(Lit+u^3)^{4/3}}{4Li/3}]_0^t"


"\Rightarrow x=\frac{3}{4Li}[(Lit+u^3)^{4/3}-u^4]"


"\Rightarrow4Lix=3[(Lit+u^3)^{4/3}-u^4]"